OpenStudy (anonymous):

What is 0.35 written as a fraction in simplest form? A:3 and 1/2 B:3/5 C:7/20 D:1/3

OpenStudy (paki):

7/20 = 0.35

OpenStudy (texaschic101):

to do these types of problems, you need to know place value. Can you tell me what place value the 5 is in ?

OpenStudy (anonymous):

No im not sure

OpenStudy (texaschic101):

the 5 is in the hundredths place...so put 35/100, then reduce. answer is 7/20.

OpenStudy (anonymous):

Thank you

OpenStudy (whpalmer4):

0.1 = 1/10 = "one tenth" 0.01 = 1/100 = "one one-hundredth" so this is 35/100 because you need to use the 0.01 place to get the rightmost digit

OpenStudy (texaschic101):

you need to learn place values.. http://www.enchantedlearning.com/math/decimals/placevalue/

OpenStudy (texaschic101):

because if your number was .352....then your fraction would be 352/1000 because the 2 is in the thousandths place. And of course, if it reduces, then reduce it.